OpenStudy (anonymous):

the area of circle of radius r is the same as the area of a square of side a. ratio of r to a is??

OpenStudy (anonymous):

@perl

OpenStudy (chosenmatt):

lol i don't know the answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@taylor12344 @Isaiah99

OpenStudy (anonymous):

ratio of r to a = area of circle/area of square

OpenStudy (anonymous):

\[area of circle = \pi r^2\] \[area of square = \ a^2\] \[area of circle = area of square\] \[\pi r^2 = a^2\] now find r/a

OpenStudy (anonymous):

the answer must be 1:1 a:rpi a:r

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

\[r^2/a^2 = 1/\pi \] \[r/a = \sqrt{1/\pi }\]

OpenStudy (anonymous):

ok ty @sangya21

OpenStudy (anonymous):

@satellite73